the solution to a system of two linear equations is (4,-3). one equation has a slope of 4. the slope of the other line is the negative recipical of the slope of the first. find the system of equations
Answer by mananth(16949)

(4,-3) slope = 4
m= 4

Plug value of the slope and point ( 4 , -3 ) in
Y = m x + b
-3.00 = 16 + b
b= -3 - 16
b= -19
So the equation will be
Y = 4 x -19
-4x+y=-19............................1
-----
(4,-3) slope = -1/4
m= - 1/ 4

Plug value of the slope and point ( 4 , -3 ) in
Y = m x + b
-3.00 = -1 + b
b= -3 - -1
b= -2
So the equation will be
Y = - 1/4 x -2
4y=-x-8
4y+x=-8 ........................2
m.ananth@hotmail.ca